OpenStudy (anonymous):

how to do this exercices 1+3/1+1/(1+1/3)

OpenStudy (anonymous):

just do the Division first and then do the Multiplication and then do the sum @leenaaman

OpenStudy (anonymous):

but it is a fraction inside a fraction inside a fraction @Learner11

OpenStudy (anonymous):

Let's show me my work

OpenStudy (anonymous):

1+3/1+1/(1+1/3) =1+3 +1/(1+0.33) =1+3+1/1.33 =1+3+0.75 =4.75

OpenStudy (anonymous):

@leenaaman

OpenStudy (anonymous):

Thank you @learner11

OpenStudy (anonymous):

Most welcome :D

OpenStudy (anonymous):

I'm not sure that's right - based on what the poster said, I believe they're trying to evaluate \[ 1 + \frac{3}{1 + \frac{1}{1+\frac{1}{3}}}\] In which case the evaluation proces should be \[ 1 + \frac{3}{1 + \frac{1}{1+\frac{1}{3}}}\] \[ =1 + \frac{3}{1 + \frac{1}{4/3}}\] \[= 1 + \frac{3}{1 + \frac{3}{4}}\] \[= 1 + \frac{3}{7/4}\] \[= 1 + \frac{12}{7}\] \[= \frac{19}{7}\]